OpenStudy (anonymous):

A customer sits in an amusement park ride in which the compartment is to be pulled downwards in the negative direction of y axis with accelaration magnitude of 1.24 g(g=9.8 m/sec ^2). A 0.567 gms coin is placed on customers knee. once motion begins, find coins accelaration relative to a)ground b)customer c) how long does the coin take to reach the compartment ceiling 2.2 m above the knee. d)the actual force on the coin e)the apparent force according to customer's measure of coins accelaration

OpenStudy (anonymous):

keep in mind the properties of frame and solve these by following way: a)acceleration=1.24*g-g=0.24*g upward b)\[a(coin,costumer)=a(coin,ground)-a(customer,ground)=0.24-0.24=0\] c)apply \[s=u \times t +(1\div2)a \times t ^{2}\] here u=0;and s=2.2m and a=0.24*g....find t. d) actual force on the coin=m*0.24g e)apparent force according to costumer=m*g

OpenStudy (anonymous):